OpenStudy (anonymous):

What are the exact solutions of x2 = 5x + 2? x = x equals 5 plus or minus the square root of thirty-three all over 2 x = x equals negative 5 plus or minus the square root of thirty-three all over 2 x = x equals 5 plus or minus the square root of seventeen all over 2 x = x equals negative 5 plus or minus the square root of seventeen all over 2

OpenStudy (nurali):

x ² - 5x - 2 = 0 x = [ - b ± √ ( b ² - 4 a c ) ] / 2 a x = [ 5 ± √ ( 25 + 8 ) ] / 2 x = [ 5 ± √ ( 33 ) ] / 2

OpenStudy (anonymous):

so it wht? a b c d

OpenStudy (nurali):

x = x equals negative 5 plus or minus the square root of thirty-three all over 2

OpenStudy (anonymous):

oh okY Thx

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

its not negative 5 because it is -b in the formula which means -(-5) which is positive 5

OpenStudy (anonymous):

x = x equals 5 plus or minus the sqr root of 33 all over 2